OpenStudy (anonymous):

work that problem please medal given 5/2+x + x/x-4

OpenStudy (anonymous):

x^2 + 7x - 20 ------------ x^2 - 2x - 2

OpenStudy (anonymous):

I didn't want to write it all out so I used a website to type it for me

OpenStudy (anonymous):

And then at the end, you multiply both sets at the bottom together and then you are finished. This website doesn't do that part for some reason.

OpenStudy (anonymous):

OMGGGGGG THANK YOU CAN YO HELP WITH ONE MORE ?

OpenStudy (anonymous):

Sure :)

OpenStudy (anonymous):

9/3x-1 - 5x/2x+3

OpenStudy (anonymous):

9 5x ---- - ----- 3x-1 2x+3 like that?

OpenStudy (anonymous):

yessss ( : perfect

OpenStudy (anonymous):

omggg thank you what website is that ?

OpenStudy (anonymous):

Your COMPLETE answer is like this: 15x^2 - 23 -27 ------------- 6x^2 + 15x + 9 http://www.wolframalpha.com/input/?i=simplify+%289%29%2F%283x-1%29+-+%285x%29%2F%282x%2B3%29 Hint: if there are three answers to your problem under the "Results:" tab, the middle one is your best shot kay darlin!

OpenStudy (anonymous):

okay lolll thank you i use wolfram all the time but that didnt pop up lolll

OpenStudy (anonymous):

hahaha! yeah it's awesome

OpenStudy (anonymous):

@RICARDOismyfirstname i just typed this in on wolfram you said the middle one so do this look right ? simplify 1/2+3/t=5/8

OpenStudy (anonymous):

also how do you find out how wolfram show the work ?

OpenStudy (anonymous):

You go to "step-by-step solutions"

OpenStudy (anonymous):

im sorry where do i find it lolll im lost : (

OpenStudy (anonymous):

okay so check this. On the bottom right of the first picture there is a tab that says, "Step-by-step solutions" it is also on the top right of the second picture.

OpenStudy (anonymous):

okay i gotchu ( : so where do i find that when i pu in this ?simplify 1/2+3/t=5/8

OpenStudy (anonymous):

you think you could write that out for me?

OpenStudy (anonymous):

omggg i think you have to sign in to get wolfram pro for step by step soultions i got it thank youuuu

OpenStudy (anonymous):

oh okay:) and if you just hit the "x" when that "wolfram alpha pro" pops up, just hit it and it will still let you get it.

OpenStudy (anonymous):

ohhhh okay without paying for it, wowwww thank you so muchhhhh you were a big help!!!

OpenStudy (anonymous):

No prob doll! Have a good one!

OpenStudy (anonymous):

you tooo ( :